# Beyond Needle(s) in the Embodied Haystack: Environment, Architecture, and Training Considerations for Long Context Reasoning

### v1

@@ 1 Introduction
We introduce $\infty$-Thor, a new framework for generation, training, and evaluation of long-horizon embodied tasks. Our benchmark uniquely features tasks with a synthetic final goal, which involves multiple objects that appear at distant time steps, requiring multi-step reasoning across over hundreds of steps. Figure 1 illustrates an example: the agent observes the tomato at an early step (t=17) and the counter top much later (t=560). Then, the final task is given at t=670, which requires the agent to place the tomato on the counter top. This setup highlights the challenge of long-horizon dependency, where key objects and locations must be remembered and acted upon after hundreds of steps.
This long-horizon setup introduces a new challenging task, Needle(s) in the Embodied Haystack (NiEH). Unlike the standard Needle in a Haystack task liu-etal-2024-lost , which focuses on recalling a single clue in text, NiEH poses two main challenges: (1) multiple scattered clues (Needles) and (2) multi-modal inputs that combine visual and linguistic observations from the environment (Embodiment). This task is designed to evaluate the agent’s ability to recall and reason about previously encountered environmental details, such as identifying objects and recalling performed actions.
@@
We further investigate various architectural considerations for embodied agents to operate under extreme sequence lengths. We show that interleaved Goal-State-Action modeling—a multimodal, goal-conditioned VLA architecture that jointly models interleaved sequences of goals, states, and actions using a LLM backbone is the most practical approach for this class of problems. Moreover, since standard LLMs are constrained by fixed context windows and cannot natively handle inputs exceeding 1M tokens, we explore long-context extension techniques such as rotary embedding scaling and positional interpolation chen2023extendingcontextwindowlarge ; ding2024longropeextendingllmcontext ; peng2024yarn . Lastly, we demonstrate how to further strengthen long-context reasoning by fine-tuning the model on extended-context inputs using Context Parallelism, a parallel training strategy that allows efficient scaling to very long sequences.
@@ 2 Related Work
Table 1: Comparison of benchmarks. We use Short (< 50 steps), Medium (50–300 steps), and Long (> 300 steps) to describe task horizon, reflecting the approximate number of environment steps required to complete a task in each benchmark. Single/Multi in the QA set column denotes single- and multi-evidence question type.
Benchmark Task Interaction Dataset QA set
/Platform Horizon w/ env modality # steps GT actions single multi
ProcTHOR deitke2022procthorlargescaleembodiedai
MineDojo fan2022minedojobuildingopenendedembodied Long
Habitat 3.0 puig2023habitat30cohabitathumans Long
VirtualHome puig2018virtualhomesimulatinghouseholdactivities Short multi 11.6
ALFRED shridhar2020alfred Medium multi 50
ALFWorld ALFWorld20 Medium text 50
BEHAVIOR-100 srivastava2021behaviorbenchmarkeverydayhousehold Med/Long
BALROG author2024BALROG Long
EQA embodiedqa
MM-EGO ye2025mmego
$\infty$-Thor Long multi 600+
@@
Long-horizon Planning in Virtual Environments. AI2THOR kolve2017ai2 provides interactive indoor environments widely used for embodied reasoning research, while ProcTHOR deitke2022procthorlargescaleembodiedai extends these capabilities by procedurally generating scalable environments, potentially facilitating longer trajectories. MineDojo fan2022minedojobuildingopenendedembodied offers an open-ended platform within Minecraft, explicitly geared toward tasks requiring extensive long-term planning. Additionally, platforms such as VirtualHome puig2018virtualhomesimulatinghouseholdactivities and Habitat 3.0 puig2023habitat30cohabitathumans have demonstrated suitability for tasks involving long-term interactions and complex activity sequences. However, all of these platforms only provide environments and do not include standardized datasets or benchmark suites to support training and evaluation for long-horizon embodied tasks.
@@
Embodied QA and Multimodal Needle in the Haystack Tasks. Embodied QA tasks, such as EmbodiedQA embodiedqa and MM-EGO ye2025mmego , require agents to answer questions grounded in visual observations, often demanding spatial and temporal reasoning over the agent’s trajectory. While these benchmarks emphasize multimodal understanding, they do not involve active interaction with the environment during evaluation. Another related area to our NiEH task is the multimodal Needle in a Haystack (NiH) problem. Traditional NiH tasks primarily assessed textual recall within long-context inputs liu-etal-2024-lost , while recent multimodal variants extend this idea by incorporating visual components wang2024needlemultimodalhaystack ; wang-etal-2025-multimodal . However, these works operate on relatively short context windows (typically up to 72K tokens) and do not require embodied reasoning or temporal dependencies across a dynamic trajectory.
@@
Datasets and Benchmarks for Long-horizon Embodied Tasks. Recent efforts have pushed toward long-horizon embodied tasks, where agents must complete multi-step goals with extended temporal dependencies. While ALFRED shridhar2020alfred and ALFWorld ALFWorld20 introduced multi-step instruction-following tasks with action annotations and textual grounding, their task horizons are relatively short, typically under 50 steps. BEHAVIOR-100 srivastava2021behaviorbenchmarkeverydayhousehold evaluates agent generalization on household activities, some of which require prolonged engagement, but mainly focus on single task. BALROG author2024BALROG is a benchmark for testing the agentic capabilities of long-context LLMs, but its scope is limited to games.
@@
Long-context Benchmarks. Outside embodied AI, general benchmarks have addressed challenges in long-context reasoning. Benchmarks, such as LongBench bai-etal-2024-longbench and RULER hsieh2024rulerwhatsrealcontext , focus on retrieval or summarization tasks. GSM-$\infty$ zhou2025gsminfinitellmsbehaveinfinitely extends GSM-8K cobbe2021trainingverifierssolvemath to assess mathematical reasoning over extremely long textual inputs.
@@ 3 $\infty$-Thor: An Environment for Generating, Training, and Evaluating Long-horizon Embodied Tasks
$\infty$-Thor features with a generation framework for synthesizing long trajectories to train and evaluate AI agents in long-horizon embodied tasks. We build $\infty$-Thor upon the AI2-THOR simulator kolve2017ai2 , an interactive 3D environment for embodied AI research that supports diverse scenes, objects, and agent actions. $\infty$-Thor enables the creation of trajectories with no length limit, and provides an evaluation setup where agents can interact dynamically with the environment during both training and testing. This supports both offline learning by producing large-scale datasets, and online learning through direct agent-environment interaction.
Each trajectory generated by $\infty$-Thor consists of multiple task goals, such as “Put a clean sponge on a metal rack” and “Pick up the apple and place it on the microwave”, requiring grounded understanding and action to achieve the goal. This enables the agent to explore and interact with the environment over extended episodes. At the end of each trajectory, the agent is assigned a synthetic long-horizon task that requires reasoning over entities encountered at distant time steps. For the example in Figure 1, the long-horizon task (Sub-goal #23) at step t=689, “Put the tomato on the counter top”, depends on observations made far earlier: the tomato at t=17 and the counter top at t=560. Our generation framework can generate unlimited tasks, the trajectories can be exceptionally long, exceeding 1M context tokens or beyond when the trajectory is processed with LLMs.
Successfully completing this task requires the agent to (1) memorize and integrate key environmental information over hundreds of steps, and (2) plan actions based on dependencies that are separated in time, demonstrating the need for long-context reasoning and robust spatio-temporal memory.
@@ 3.1 Static Evaluation: Needle(s) in the Embodied Haystack
After generating QA pairs and annotating the GT steps, we cross-validate the answerability of each question with GT images using four different multimodal LLMs: LLaVA-OneVision 7B li2024llava , Qwen2.5-VL 7B bai2025qwen25vltechnicalreport , Deepseek-VL 7B lu2024deepseekvlrealworldvisionlanguageunderstanding , and Pixtral 12B agrawal2024pixtral12b . Since these models are highly capable at standard visual QA, we filter out the questions that none of the four models successfully answer with GT images. At test time, the entire trajectory is treated as a Haystack and then cropped based on the GT image’s depth. Full details on templates, generation rules, and the validation scores of the four models are included in the Appendix.
@@
Table 2: Dataset Statistics
@@ 3.2 Constructing Long-horizon Trajectories for Interactive Evaluations
With $\infty$-Thor’s generation framework, we can synthesize long-horizon trajectories to construct training, validation, and test sets for offline learning and evaluation. Our approach builds upon a planner-based method kolve2017ai2 , in which we sequentially concatenate multiple single-task demonstrations into a extended trajectory, while maintaining consistency in object states and agent interactions throughout.
To generate each trajectory, we first sample a task type from one of seven predefined templates (e.g., pick two objects and place, pick and place with movable receptacle). We then sample objects that are required to perform the task, such as items to be picked up or receptacles to be interacted with. Based on the sampled task and objects, we use a classical task planner that operates on PDDL-defined domains to generate ground-truth action sequences. These plans are executed in a simulator, and only successful rollouts are retained. We then concatenate these successful demonstrations to construct long-horizon sequences that span hundreds of steps. For the final goal, the involved objects are sampled exclusively from those seen during the early 20% and the final 20% of the trajectory. This enforces a long-term temporal dependency between two objects that must be jointly referenced to complete the final task. Through this procedure, we generate 2,456/125/225 trajectories for the training, validation, and test sets, respectively. Details on task types and a pseudo-algorithm for the generation process are provided in the Appendix.
@@ 4 Architectures for Long-Horizon Vision-Language-Action Models
Embodied agents must effectively interact with complex, dynamic environments, necessitating capabilities to interpret multimodal inputs (vision, language) and produce coherent sequences of actions. Developing such Vision-Language-Action (VLA) models is particularly challenging due to the need for seamless integration of perceptual understanding, linguistic reasoning, and action prediction. Existing VLA models either use separate encoders for vision, language, and action modules shridhar2020alfred , or focus on short-horizon, low-level tasks such as robot arm manipulation in constrained environments brohan2023rt1roboticstransformerrealworld ; kim24openvla , where decisions depend only on the most recent observation and a single instruction. While recent multimodal LLMs like LLaVA NEURIPS2023_6dcf277e , MiniGPT-4 zhu2023minigpt , and Llama 3.2 meta2024llama32 show strong multimodal reasoning abilities, they primarily operate on static inputs (e.g., single or few images) and lack the dynamic interactivity and memory needed for long-horizon embodied tasks involving continuous vision-language-action sequences. Moreover, many state-of-the-art models are only accessible via proprietary APIs, making them impractical for real-time, controllable embodied settings and managing long-term memory states.
@@
Context Extension. Given the limitations in context length of most LLMs, using off-the-shelf models is insufficient for processing long inputs such as those exceeding 1M tokens. We explore various long-context extension techniques that allows the model to generalize to longer input sequences without retraining from scratch. Specifically, we consider: Linear Interpolationchen2023extendingcontextwindowlarge : Rescales input positions to fit within the pretrained RoPE range by linearly interpolating positional indices; Dynamic Scalingchen2023extendingcontextwindowlarge : Adapts RoPE frequencies at runtime based on the input sequence length, using a linear rescaling to maintain consistent positional encoding behavior across varying lengths; YaRNpeng2024yarn : Dynamically interpolates attention frequencies during inference, balancing between pretrained and extrapolated positional regimes; LongRoPEding2024longropeextendingllmcontext : Augments RoPE with specially designed extrapolation functions, enabling robust generalization to long sequences without degrading attention quality. We apply these techniques during fine-tuning, at inference time, or both.
@@
Context Parallelism leverages Ring Attention liu2023ring , a novel parallel implementation of the attention layer. In Ring Attention, key-value (KV) shards are cyclically shuffled across devices, and partial attention scores are computed iteratively. This process is repeated until all KV shards have been incorporated on each device, ensuring complete attention coverage without the full memory cost of standard attention. By combining Context Parallelism with our dataset of extended long-context inputs, we are able to scale fine-tuning to substantially longer sequences, unlocking improved long-horizon reasoning capabilities.
@@ 5.1 Static Evaluation: Needle(s) in the Embodied Haystack
Results. Figure 4 presents the performance of LLaVA-OneVision 7B li2024llava model across various context extension methods. Linear Interpolation, Dynamic Scaling, and LongRoPE scaling all struggled with very long contexts beyond 128K tokens (the results of Linear Interpolation are excluded from Figure since it fails at all examples). YaRN consistently outperformed other methods across both NiEH and NsiEH, successfully answering questions at context lengths exceeding 384K tokens, likely due to its architectural alignment with LLaVA-OneVision’s Qwen2 LM backbone, which employs RoPE and YaRN scaling during pretraining. YaRN performed best at moderate scaling factors (e.g., x4), however, further scaling to x8 and x16 did not yield additional gains. In particular, x16 slightly improved performance in the 256K–384K token range but led to degradation notably at shorter context sizes (<64K), suggesting that excessive scaling may introduce instability and negatively impact performance. Overall, all methods fail beyond 512K tokens. We expect these trends to persist as models are pushed to even larger context windows, highlighting the need for more advanced methods to effectively handle extremely long-context scenarios in the future.
@@
Figure 4: Results of Needle(s) in the Embodied Haystack with the LLaVA-OneVision 7B model. The white dashed line denotes the maximum input context length of the model. Context Parallelism is applied to all experiments with the context size over 384K.

### v2

@@ 1 Introduction
We introduce $\infty$-Thor, a new framework for generation, training, and evaluation of long-horizon embodied tasks. Our benchmark uniquely features tasks with a synthetic final goal, which involves multiple objects that appear at distant time steps, requiring multi-step reasoning across over hundreds of steps. Figure 1 illustrates an example: the agent observes the tomato at an early step (t=17) and the counter top much later (t=560). Then, the final task is given at t=670, which requires the agent to place the tomato on the counter top. This setup highlights the challenge of long-horizon dependency, where key objects and locations must be remembered and acted upon after hundreds of steps. Beyond these long-horizon dependencies, our framework also generates low-level robot-arm manipulation action sequences aligned with the trajectories, enabling agents to bridge from high-level reasoning to fine-grained physical execution.
This long-horizon setup introduces a new challenging task, Needle(s) in the Embodied Haystack (NiEH). Unlike the standard Needle in a Haystack task (Liu et al., 2024), which focuses on recalling a single clue in text, NiEH poses two main challenges: (1) multiple scattered clues (Needles) and (2) multi-modal inputs that combine visual and linguistic observations from the environment (Embodiment). This task is designed to evaluate the agent’s ability to recall and reason about previously encountered environmental details, such as identifying objects and recalling performed actions.
@@
We further investigate various architectural considerations for embodied agents to operate under extreme sequence lengths. We show that interleaved Goal-State-Action modeling—a multimodal, goal-conditioned VLA architecture that jointly models interleaved sequences of goals, states, and actions using a LLM backbone is the most practical approach for this class of problems. Moreover, since standard LLMs are constrained by fixed context windows and cannot natively handle inputs exceeding 1M tokens, we explore long-context extension techniques such as rotary embedding scaling and positional interpolation (Chen et al., 2023; Ding et al., 2024; Peng et al., 2024). Lastly, we demonstrate how to further strengthen long-context reasoning by fine-tuning the model on extended-context inputs using Context Parallelism, a parallel training strategy that allows efficient scaling to very long sequences.
@@ 2 Related Work
Table 1: Comparison of benchmarks. We use Short (<50<50 steps), Medium (5050-300300 steps), and Long (>300>300 steps) to describe task horizon, reflecting the approximate number of environment steps required to complete a task in each benchmark. (Inter. w/ env: Interaction with the environment, Mod: Modality, GT: GT actions; single/multi in the QA set column denotes single- and multi-evidence question type. * indicates the number of annotations newly collected in that work.)
Benchmark / Platform Task Inter. Dataset QA set
Horizon w/ env Mod Avg steps GT single multi
ProcTHOR (Deitke et al., 2022)
MineDojo (Fan et al., 2022) Long
Habitat 3.0 (Puig et al., 2023) Long
VirtualHome (Puig et al., 2018) Short multi 11.6
ALFRED (Shridhar et al., 2020) Medium multi 50
ALFWorld (Shridhar et al., 2021) Medium text 50
BEHAVIOR-100 (Srivastava et al., 2021) Med/Long
BALROG (Paglieri et al., 2024) Long
EAI (Li et al., 2024b) Med/Long 14.6*
EQA (Das et al., 2018)
MM-EGO (Ye et al., 2025)
$\infty$-Thor $\infty$ multi 627
@@
Long-horizon Planning in Virtual Environments. AI2THOR (Kolve et al., 2017) provides interactive indoor environments widely used for embodied reasoning research, while ProcTHOR (Deitke et al., 2022) extends these capabilities by procedurally generating scalable environments, potentially facilitating longer trajectories. MineDojo (Fan et al., 2022) offers an open-ended platform within Minecraft, explicitly geared toward tasks requiring extensive long-term planning. Additionally, platforms such as VirtualHome (Puig et al., 2018) and Habitat 3.0 (Puig et al., 2023) have demonstrated suitability for tasks involving long-term interactions and complex activity sequences. However, all of these platforms only provide environments and do not include standardized datasets or benchmark suites to support training and evaluation for long-horizon embodied tasks.
@@
Embodied QA and Multimodal Needle in the Haystack Tasks. Embodied QA tasks, such as EmbodiedQA (Das et al., 2018) and MM-EGO (Ye et al., 2025), require agents to answer questions grounded in visual observations with spatial and temporal reasoning, but without active environment interaction during evaluation. Our NiEH task is also related to multimodal Needle in a Haystack (NiH) problems. While early NiH focused on textual recall in long contexts (Liu et al., 2024), recent multimodal extensions add visual inputs (Wang et al., 2024; 2025), though they remain limited to shorter contexts (up to 72K tokens) and lack embodied reasoning or temporal dependencies.
@@
Datasets and Benchmarks for Long-horizon Embodied Tasks. Recent efforts have pushed toward long-horizon embodied tasks, where agents must complete multi-step goals with extended temporal dependencies. While ALFRED (Shridhar et al., 2020) and ALFWorld (Shridhar et al., 2021) introduced multi-step instruction-following tasks with action annotations and textual grounding, their task horizons are relatively short, typically under 50 steps. BEHAVIOR-100 (Srivastava et al., 2021) evaluates agent generalization on household activities, some of which require prolonged engagement, but mainly focus on single task. BALROG (Paglieri et al., 2024) is a benchmark for testing the agentic capabilities of long-context LLMs, but its scope is limited to games. Recently, EAI (Li et al., 2024b) proposed a generalized interface to evaluate LLMs for embodied decision making, while our framework focuses on multimodal online evaluation with real-time rewards and a new long-horizon reasoning task (NiEH).
@@
Long-context Benchmarks. Outside embodied AI, general benchmarks have addressed challenges in long-context reasoning. Benchmarks, such as LongBench (Bai et al., 2024) and RULER (Hsieh et al., 2024), focus on retrieval or summarization tasks. GSM-$\infty$ (Zhou et al., 2025) extends GSM-8K (Cobbe et al., 2021) to assess mathematical reasoning over extremely long textual inputs. More recently, LMAct (Ruoss et al., 2025) proposed a benchmark for evaluating frontier models’ long-context multimodal decision-making on interactive game-based tasks, with up to 1M context lengths.
@@ 3 $\infty$-Thor: An Environment for Generating, Training, and Evaluating Long-horizon Embodied Tasks
$\infty$-Thor features with a generation framework for synthesizing long trajectories to train and evaluate AI agents in long-horizon embodied tasks. We build $\infty$-Thor upon AI2-THOR (Kolve et al., 2017) simulator, an interactive 3D environment for embodied AI research that supports diverse scenes, objects, and agent actions. $\infty$-Thor enables the creation of unlimited trajectories with arbitrarily long, and provides an evaluation setup where agents can interact dynamically with the environment during both training and testing. This supports both offline learning by producing large-scale datasets, and online learning through direct agent-environment interaction.
Each trajectory generated by $\infty$-Thor consists of multiple task goals, such as “Put a clean sponge on a metal rack” and “Pick up the apple and place it on the microwave”, requiring grounded understanding and action to achieve the goal. At the end of each trajectory, the agent is assigned a synthetic long-horizon task that requires reasoning over entities encountered at distant time steps. For the example in Figure 1, the long-horizon task (Sub-goal #23) at step t=689, “Put the tomato on the counter top”, depends on observations made far earlier: the tomato at t=17 and the counter top at t=560.
Our generation framework can generate unlimited tasks, the trajectories can be exceptionally long, exceeding 1M context tokens or beyond when the trajectory is processed with LLMs. Moreover, $\infty$-Thor can generate low-level robot manipulation trajectories which are compatible with the ManipulaTHOR (Ehsani et al., 2021) simulator, supporting both symbolic plans and low-level controls. Successfully completing this task requires the agent to (1) memorize and integrate key environmental information over hundreds of steps, and (2) plan actions based on dependencies that are separated in time, demonstrating the need for long-context reasoning and robust spatio-temporal memory, and (3) execute low-level physical actions through extended manipulation sequences.
@@ 3.1 Static Evaluation: Needle(s) in the Embodied Haystack
After generating QA pairs and annotating the GT steps, we cross-validate the answerability of each question with GT images using four different multimodal LLMs: LLaVA-OneVision 7B (Li et al., 2024a), Qwen2.5-VL 7B (Bai et al., 2025), Deepseek-VL 7B (Lu et al., 2024), and Pixtral 12B (Agrawal et al., 2024). Since these models are highly capable at standard visual QA, we filter out the questions that none of the four models successfully answer with GT images. At test time, the entire trajectory is treated as a Haystack and then cropped based on the GT image’s depth. Full details on templates, generation rules, and the validation scores of the four models are included in the Appendix.
@@ 3.2 Constructing Long-horizon Trajectories for Interactive Evaluations
With $\infty$-Thor’s generation framework, we can synthesize long-horizon trajectories to construct training, validation, and test sets for offline learning and evaluation. Our approach builds upon a planner-based method (Kolve et al., 2017), in which we sequentially concatenate multiple single-task demonstrations into a extended trajectory, while maintaining consistency in object states and agent interactions throughout.
To generate each trajectory, we first sample a task type from one of seven predefined templates (e.g., pick two objects and place, pick and place with movable receptacle). We then sample objects that are required to perform the task, such as items to be picked up or receptacles to be interacted with. Based on the sampled task and objects, we use a classical task planner that operates on PDDL-defined domains to generate ground-truth action sequences. Only successful rollouts (re-simulated without failure) are retained, ensuring the reproducibility and reliability of our trajectories. We then concatenate these successful demonstrations to construct long-horizon sequences that span hundreds of steps.
Table 2: Dataset Statistics. “low” denotes low-level robot arm manipulation actions.
@@
# avg steps (low) 3,000+ 5,000+ 5,000+
For the final goal, the involved objects are sampled exclusively from those seen during the early 20% and the final 20% of the trajectory. This enforces a long-term temporal dependency between two objects that must be jointly referenced to complete the final task. Through this procedure, we generate 2,456/125/225 trajectories for the training, validation, and test sets, respectively.
We use a similar approach for low-level manipulation actions. For the Pick-up action, a trajectory is generated from the robot arm’s start point to the object’s target point (reversed for the Put action). Each trajectory is executed through fine-grained arm movements $(\Delta x,\Delta y,\Delta z)$, and only successful rollouts are retained. Details on task types and a pseudo-algorithm for the generation process are provided in the Appendix.
@@ 4 Architectures for Long-Horizon Vision-Language-Action Models
Embodied agents must effectively interact with complex, dynamic environments, necessitating capabilities to interpret multimodal inputs (vision, language) and produce coherent sequences of actions. Developing such Vision-Language-Action (VLA) models is particularly challenging due to the need for seamless integration of perceptual understanding, linguistic reasoning, and action prediction. Existing VLA models either use separate encoders for vision, language, and action modules (Shridhar et al., 2020), or focus on short-horizon in constrained environments (Brohan et al., 2023; Kim et al., 2024), where decisions depend only on the most recent observation and a single instruction. While recent multimodal LLMs like LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023), and Llama 3.2 (AI, 2024) show strong multimodal reasoning abilities, they primarily operate on static inputs (e.g., single or few images) and lack the dynamic interactivity and memory needed for long-horizon embodied tasks involving continuous vision-language-action sequences. Moreover, many state-of-the-art models are only accessible via proprietary APIs, making them impractical for real-time, controllable embodied settings and managing long-term memory states.
@@
Context Extension. Given the limitations in context length of most LLMs, using off-the-shelf models is insufficient for processing long inputs such as those exceeding 1M tokens. We explore various long-context extension techniques that allows the model to generalize to longer input sequences without retraining from scratch. Specifically, we consider: Linear Interpolation(Chen et al., 2023): Rescales input positions to fit within the pretrained RoPE range by linearly interpolating positional indices; Dynamic Scaling(Chen et al., 2023): Adapts RoPE frequencies at runtime based on the input sequence length, using a linear rescaling to maintain consistent positional encoding behavior across varying lengths; YaRN(Peng et al., 2024): Dynamically interpolates attention frequencies during inference, balancing between pretrained and extrapolated positional regimes; LongRoPE(Ding et al., 2024): Augments RoPE with specially designed extrapolation functions, enabling robust generalization to long sequences without degrading attention quality. We apply these techniques during fine-tuning, at inference time, or both.
@@
Context Parallelism leverages Ring Attention (Liu et al., 2023a), a novel parallel implementation of the attention layer. In Ring Attention, key-value (KV) shards are cyclically shuffled across devices, and partial attention scores are computed iteratively. This process is repeated until all KV shards have been incorporated on each device, ensuring complete attention coverage without the full memory cost of standard attention. By combining Context Parallelism with our dataset of extended long-context inputs, we are able to scale fine-tuning to substantially longer sequences, unlocking improved long-horizon reasoning capabilities.
@@ 5.1 Static Evaluation: Needle(s) in the Embodied Haystack
Figure 4: Results of Needle(s) in the Embodied Haystack. The white dashed line denotes the maximum input context length of the model. Qwen2.5-VL was pre-trained initially with an 8K token context window and incrementally scaled up to 32K tokens in subsequent stages (Bai et al., 2025). The gray area indicates contexts not applicable (N/A) due to the model’s smaller image token size limiting sequences to under 128K tokens. Context Parallelism is applied to all experiments with the context size over 384K.
Results. Figure 4 presents the performance of LLaVA-OneVision 7B (Li et al., 2024a) model across various context extension methods. Linear Interpolation, Dynamic Scaling, and LongRoPE scaling all struggled with very long contexts beyond 128K tokens (the results of Linear Interpolation are excluded from Figure since it fails at all examples). YaRN consistently outperformed other methods, effectively handling contexts above 384K tokens, likely due to its architectural alignment with LLaVA-OneVision’s Qwen2 LM backbone, which employs RoPE and YaRN scaling during pretraining. YaRN performed best at moderate scaling factors (e.g., x4), however, further scaling to x8 and x16 did not yield additional gains. In particular, x16 slightly improved performance in the 256K–384K token range but led to degradation notably at shorter context sizes (¡64K), suggesting that excessive scaling may introduce instability and negatively impact performance. Overall, all methods fail beyond 512K tokens, highlighting a need for improved long-context methods.
Figure 4 also includes additional results from DeepSeek-VL 7B (Lu et al., 2024), Qwen2.5-VL 7B (Bai et al., 2025), and Gemini 2.0 Flash [^1] We used the gemini-2.0-flash-001 version for all experiments. (Google DeepMind, 2024). Each model processes images into tokens differently—DeepSeek (576 tokens/image), Qwen2.5 (121 tokens/image), and Gemini 2.0 Flash (258 tokens/image)—which consequently impacts maximum context lengths when transforming N(s)iEH sequences into tokenized inputs (DeepSeek-VL: 512K, Qwen2.5-VL: 128K, Gemini 2.0 Flash: 256K). DeepSeek-VL struggles beyond its 4K pretrained context limit. Qwen2.5-VL maintains stronger performance up to approximately 64K tokens, benefiting from its smaller per-image tokenization, but performance notably declines on NsiEH at longer contexts. Gemini 2.0 Flash performs robustly up to 8K tokens but degrades beyond 128K, especially on NsiEH tasks, highlighting room for improvement in complex long-range multimodal reasoning.